The area of a sail is 40 1/2 square feet. The base and the height of the sail are equal. What is the height of the sail (in feet).
I tried 40 1/2=1/2 x 2x
Thank you for your help.

Answer by JulietG(1812)

That's an odd value.
.
H * W = 40.5
Since they are equal,
H^2 = 40.5
Take the square root of each:
H = 6.364
(and therefore W = 6.364)
.
If the value is something else, simply take its square root.
OR, possibly...
Is the sail triangular rather than square? If so, then solve it this way:
h^2 = 1/2 40.5
h^2 = 20.25
h = 4.5 feet
That makes a lot more sense. This question really should be clarified as there are different sorts of "sails" and rarely would one be a triangle yet have the same base and height. *sigh*